OpenStudy (anonymous):

Which of the following are solutions to the equation? (check all that apply) a. -6 b. -5 c.6 d.4 e.-2 f.5

OpenStudy (anonymous):

whoops the equation is x2-3x+27=8x-3

OpenStudy (anonymous):

is that x squared or 2x?

OpenStudy (anonymous):

x squared

OpenStudy (anonymous):

ok so first you move the 8x to the other side and it becomes negative, so when you combine -8x-3x=-11x. do the same with the negative 3 and it becomes positive, so you get 27+3=30.

OpenStudy (anonymous):

The next step is to factor. so you are left with x^2-11x+30=0. when you factor, you get (x-6) (x-5)=0 so x-6=0 and x-5=0. by solving these equations the answers are 6 and 5

OpenStudy (anonymous):

thank you so much!